If a bag weighs 10kg and the coefficient against the ground is 0.5.The angle between the rope and ground is 20. How much extra force is needed to pull this bag behind himself


1
Expert's answer
2021-10-20T10:41:18-0400

Let us find the force required to pull the bag at constant velocity. Use Newton's second law:


"y:-mg+N+F\sin\theta=0,\\
x: -\mu N+F\cos\theta=0,\\\space\\
N=mg-F\sin\theta,\\
F\cos\theta=\mu mg
-\mu F\sin\theta,\\\space\\
F=\frac{\mu mg}{\cos\theta+\mu\sin\theta}=44.1\text{ N}."